OpenStudy (anonymous):

one card is selected at random from a deck of cards. Determine the probability that the card selected is 22. a diamond

OpenStudy (anonymous):

can u explain how to get the answer?

OpenStudy (anonymous):

there are 13 diamonds in a pack of 52 cards, so the probablity to pull a diamond out of the pack is 13/52=1/4

OpenStudy (anonymous):

so how did u get 1/4? take and divided 13 into 52?

OpenStudy (anonymous):

yes, that's right.

jimthompson5910 (jim_thompson5910):

There are 13 diamond cards out of 52 cards total. So the probability is 13/52 = 1/4

OpenStudy (anonymous):

the other way to think about it is that there are four suits equally divided. so it is 1/4

jimthompson5910 (jim_thompson5910):

Each suit has the following: 2, 3, 4, 5, 6, 7, 8, 9, 10, J, Q, K, A (13 cards in each suit)

OpenStudy (anonymous):

ok i know that..

OpenStudy (anonymous):

ty